OpenStudy (anonymous):

Which expression is the simplified form of the square root of 605w^24 ? 12w^12 times the square root of 5 11w^12 times the square root of 5 5w^4 times the square root of 11 5w^4 times the square root of 12

OpenStudy (hoblos):

11w^12 times the square root of 5

OpenStudy (anonymous):

to show:\[=\sqrt{605w^{24}} =\sqrt{121(5)w ^{24}} = 11w ^{12} \sqrt{5}\]